OpenStudy (anonymous):

I have 4 questions! What is the value of a in the equation a = 2 + 3a + 8? 10 5 −5 −10

OpenStudy (anonymous):

please just the answers!

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

ok i have 3 others

OpenStudy (anonymous):

What is the value of y in the equation 2(3y + 6 + 3) = 196 − 16? 27 29 30 33

OpenStudy (anonymous):

27

OpenStudy (anonymous):

How many solutions does the equation 4x + 8 = −2 − 2 + 5x have? Two One None Infinitely many

OpenStudy (anonymous):

1 I think

OpenStudy (anonymous):

What is the solution to the equation 5x + 2(x − 4) = 5x + x − 10? 3 2 −2 −3 Thank you

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

omg you are amazing! i fanned metal and gave best responce!